OpenStudy (anikate):

Find the mid-point between (3, -1) and (0, 2)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\huge \color{green}{ Midpoint (x, y) = (\frac{x_1 + x_2}{2}, \frac{y _ 1 + y_2}{2})}\]

OpenStudy (anonymous):

Just add the x coordinates and divide by 2.. (Midpoint x coordinate) just add the y coordinates and divide it by 2.. (Midpoint y coordinate)

OpenStudy (anonymous):

\[x = \frac{3+0}{2}\] \[y = \frac{-1 + 2}{2}\]

OpenStudy (anonymous):

Find x and y here that will be your coordinate for Midpoint..

OpenStudy (anonymous):

@Anikate are you there??

OpenStudy (anikate):

sry my computer turned off again

OpenStudy (anikate):

thx water